OpenStudy (anonymous):

integral((x^2)/(4-3x^2)^(3/2))dx

OpenStudy (anonymous):

\[\int\limits \frac{ x^2 }{ (4-3x^2)^(3/2) }dx\]

OpenStudy (anonymous):

Trignometric substitution with a difference of squares

OpenStudy (anonymous):

\[(\sqrt{4})^2-3x^2\] \[x=\frac{ 2 }{ 3 } \sin(theta)\]

OpenStudy (gorv):

well take x^2 common from denominator

OpenStudy (anonymous):

can we use a u-sub instead of trig sub

OpenStudy (anonymous):

\[u=4-3x^2\] \[du=-6xdx\] \[dx=-(1/6x)dx\]

OpenStudy (dumbcow):

no i would stick with the trig sub, with a minor correction \[x = \frac{2}{\sqrt{3}} \sin \theta\]

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

so would i plug that into x for the whole equation including the top (x^2)

OpenStudy (dumbcow):

yes and the dx as well \[dx = \frac{2}{\sqrt{3}} \cos \theta \]

OpenStudy (dumbcow):

it should simplify to \[\frac{1}{3 \sqrt{3}} \int\limits \tan^2 \theta\]

OpenStudy (anonymous):

is that for the top part?

OpenStudy (dumbcow):

no the whole thing top part ---> sin^2 bottom part ----> cos^3 dx ----> cos tan = sin/cos

OpenStudy (anonymous):

ok, but what do you do with the rest of the equation? wolfram gives we this complex number http://www4a.wolframalpha.com/Calculate/MSP/MSP1071222f045f1f7caf0b00001g1fh4b1da7ai778?MSPStoreType=image/gif&s=12&w=443.&h=57.

OpenStudy (dumbcow):

well first do you see how it simplifies to tan^2 ?

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

ok, now we can use the identity \[\tan^2 \theta = \sec^2 \theta - 1\] this makes the integration much easier \[\frac{1}{3 \sqrt{3}} \int\limits \sec^2 \theta - 1 = \frac{1}{3 \sqrt{3}}(\tan \theta - \theta)\]

OpenStudy (dumbcow):

you follow?

OpenStudy (anonymous):

I think so hold on.

OpenStudy (dumbcow):

ok ... derivative of tan = sec^2 is a known trig identity that is why the integrating was easy

OpenStudy (anonymous):

what do you do with \[\frac{ 1 }{ 3\sqrt{3} }\tan(\theta)-\theta+c\]

OpenStudy (dumbcow):

Finally we have to plug "x" back in from original substitution \[x = \frac{2}{\sqrt{3}} \sin \theta\] \[\rightarrow \sin \theta = \frac{\sqrt{3}}{2}x\] \[\rightarrow \theta = \sin^{-1} (\frac{\sqrt{3}}{2}x)\] For the tan part, it helps to use a triangle remember SOH CAH TOA  \[\rightarrow \tan \theta = \frac{\sqrt{3} x}{\sqrt{4-3x^2}}\]

OpenStudy (anonymous):

thank you~ I got the same answer as wolfram!!!

OpenStudy (dumbcow):

yw :) sometimes you may not get exactly same as wolfram because they factor or simplify differently but it doesn't necessarily mean you are wrong